OpenStudy (anonymous):

Find the area of a trapezoid with bases of 8 and 16 and a height of 10

OpenStudy (anonymous):

do you know the equation for finding the area of a trapezoid?

OpenStudy (anonymous):

mean of (8+16)/2=12*16=192